### v1

@@ I Introduction
Prediction technology continues to advance, and multiple prediction outputs are now a staple of state-of-the-art prediction methods [tang2019multiple, choi2019drogon, phan2020covernet, salzmann2020trajectron++, shi2022motion, isele2024gaussian]. This paper examines how an autonomous driving (AD) agent can utilize multiple predictions in the behavior planning process. In the context of this work, behavior planning corresponds to the combined task of decision-making and trajectory planning.
@@
Consider the scenario depicted in Fig. 1. A pedestrian walks along a road and will likely continue straight (with 80% probability), but the pedestrian is positioned close to the street, indicating that they might turn to cross the street (with 20% probability). Selecting the most probable sequence of events results in an overly aggressive and risky behavior - we assume they will not cross and are wrong 20% of the time. However, a cautious policy brakes unnecessarily in the middle of an intersection for a pedestrian, creating confusing and potentially dangerous consequences as well. Intuitively, the solution in this instance is try to balance the two behaviors, slow down so that it is possible to brake if necessary, but not to make the overly conservative assumption that an unlikely event is true. This paper presents a formalism through which the problem of planning in the presence of multiple predictions can be expressed, where the intuitive solution of our example emerges as a result.
@@
Starting with probabilistic predictions, we can adopt a behavior that maximizes the expected reward [schmerling2018multimodal]. However, this is often insufficient for safety-critical applications because negative events such as collisions are discouraged but not forbidden. This shortcoming can be addressed by directly incorporating constraints into the formulation [ahn2021safe]. However, such an approach is still overly conservative because even though marginalizing over an agent’s distribution accurately assigns risk to both possible futures, it neglects the fact that only one future can be true, generally resulting in a behavior that avoids both possibilities.
@@
Figure 1: Illustrative example: a pedestrian at the intersection has possible intentions of walking straight or turning right. The ego vehicle (at the center of the image) must comprehend the pedestrian’s potential behaviors reflect that planning, e.g., by slowing down enough to stop safely if needed.
@@
One way around this difficulty is by delaying an agent’s decision while selecting an action that keeps a response to both futures possible [alsterda2019contingency, cui2021lookout]. In this work, we formalize the importance of delayed decision making and identify the optimal time to wait. While the generalized framework for behaving under the uncertainty associated with multiple predictions can be carried out by multiple different planning methods, to ground the formalism, we specifically detail its application to an MPC framework. The special case of solving the problem with MPC has some similarity to existing works [alsterda2019contingency, chen2022interactive]. Our approach solves parallel optimizations which was also explored by Adajania et al. [adajania2022multi], though in a different problem setting. Like our approach, Contigency MPC [alsterda2019contingency] also solves for parallel plans where the plans have different costs and constraints and a tied $0th$ state and action, however, they focus on steering under different friction assumptions, not predictions, and only issue a command for a single time-step. This avoids the decision process and the possibility of infeasibility. Branch MPC [chen2022interactive] takes into account another agent’s actions with corresponding probabilities, but unlike our work, Branch MPC reasons over the other agent’s entire action space which is substantially more computationally prohibitive, and therefore their investigation was limited to a single other agent.
@@
For an autonomous driving use case, there is a high non-convexity not addressed in many of the other works, which we show cannot be avoided in the worst case. So we present reasonable approximations that allow us to handle the non-convexities in an efficient way. As part of the approximate solution, we develop a very fast piece-wise-linear trajectory planning approximation which can enable the quick evaluation of candidate solutions, and might be useful outside of our specific setting. Finally, to demonstrate the practicality of our algorithm, as well as the realtime computation speed, we showcase our results in simulation and on 1/10 scale cars.
@@ II-B Problem Formulation
The multi-future trajectory planning (MFTP) problem is posed as a stochastic game described by the tuple $\{\mathcal{S},\mathcal{A},P,\mathcal{R}\}$, where $\mathcal{S}$ is the world state which consists of a set of states $\{S^{1},\dots,S^{n}\}$ for $n$ agents, $\mathcal{A}=\{A^{1},\dots,A^{n}\}$ is the set of actions for agents, $P:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ is the set of transition probabilities, and =$\mathcal{R}=\{R^{1},\dots,R^{n}\}$ is the set of reward functions for each agent. The ego agent’s states, actions, and reward function are represented by $S^{e}$, $A^{e}$, and $R^{e}$ respectively. The ego agent generates a trajectory $\tau=\{s_{1},\dots,s_{T}\}$ consisting of a set of states where $s_{t}\in S^{e}$ and $T$ is the time horizon.
Assumption 1
Deterministic Agent Behavior. The agents in $\mathcal{A}$ are assumed to behave deterministically, conditioned on a possible future $f\in F$.
@@
A future $f=\{s_{1}^{j},\dots,s_{T}^{j}\ ~{}\forall j\backslash{e}\}$ is defined as a set of future states for all agents starting from the current time. In practice, Assumption 1 allows us to make use of multi-intention prediction algorithms where the outputs can be taken directly from any off-the-shelf algorithm since there is no requirement that the predictions adhere to any particular constraints or have any particular distribution. Additionally, because we’re considering the trajectory as a whole, we remove the large computational burden of Branch MPC [chen2022interactive] since we no longer need to account for different actions from different agents at every time step.
@@
Note that it is sufficient to satisfy the constraints pertaining to the true future, however, that is unknown. For $i\in\{1,\dots,m\}$, we define the corresponding belief of future $f_{i}$ being $f_{*}$:
@@ III Approach
We suppose that the probability of each future trajectory is invariant for each planning instance, which allows to optimize the expected value:
$\max_{\tau}\sum_{i=1}^{m}p_{i}R(\tau|f_{i})=\max_{\tau}\mathbb{E}_{f\sim p(f)}[R(\tau|f)]\enspace.$ (3)
However, in the case of safety critical systems, some trajectories will be catastrophic. We acknowledge this by allowing for some trajectories to have a corresponding reward of $\infty$- ∞111for the case of autonomous driving, this will correspond to trajectories that result in collisions. This makes it difficult to work with expectations, since all expectations with non-zero probability of trajectories that result in $\infty$- ∞ reward will have $\infty$- ∞ expected value. We will therefore assert optimality can be sacrificed, for the purposes of minimizing catastrophic events. We will refer to this assertion as the catastrophic assertion.
@@ III-A Maximum Probability Trajectory
To ensure we have a viable action in response to an uncertain future, we want to maximize the probability that our trajectory is in the feasible set $\mathcal{T}_{*}$ of the true future $f_{*}$ at the time $t_{R}$ that the true future is revealed. We are therefore interested in maximizing the joint probability that we satisfy constraint $c_{i}$ and $\mathcal{T}_{i}$ is in $\mathcal{T}_{*}$ for each future $i$ at time $t=t_{R}$. A convenience of working with trajectories is that we know that if the ego agent is on $\tau$ at time $t_{R}$ and $\tau\in\mathcal{T}_{*}$ then it can stay on a feasible trajectory for all time after $t_{R}$.
@@
In general, it is not possible to find one single trajectory that stays in all feasible spaces, as trajectories corresponding to different futures will inevitably diverge. However we can maximize the number of constraints we are able to satisfy by keeping multiple constraints satisfied for as long as possible. Some readers might correctly identify the connection to maximum entropy here, which we will discuss in section III-D. However first we will discuss how keeping many trajectories feasible is equivalent to delaying a decision.
@@ III-B Delayed Decisions
At some time $t=t_{R}$ we will know which future is the true future $f_{*}$. Trivially this could be after time equal to the horizon time $T$ has passed, at which point the future will have already happened. In many cases however, we can observe indications or definitive decisions in advance since certain actions make subsequent future actions impossible. Note that in order to make use of any observed indications or decisions, the ego agent needs sufficient time to react.
If we know $f_{i}=f_{*}$, we can determine which states keep us in $\mathcal{T}_{i}$. Without knowing $f_{*}$ in advance, the last moment $t_{d}^{\operatornamewithlimits{e}}$ the ego agent can make a decision is defined as the point in time where there is at least one feasible trajectory that belongs to all trajectory sets $\mathcal{T}_{1},\dots,\mathcal{T}_{m}$. Note that by symmetry other agents have a similar $t_{d}^{\operatornamewithlimits{o}}$, after which point at least some futures will no longer be possible.
@@
Note that if, as our policy, we choose to make a conservative decisions immediately before $t_{d}^{\operatornamewithlimits{e}}$, we get a risk-free reduction in uncertainty (and increase in safety) for all cases where $t_{d}^{\operatornamewithlimits{o}}$ happens first. Additionally, if $t_{d}^{\operatornamewithlimits{e}}$ happens earlier, we can still benefit from waiting as a result of additional acquired information. The next section describes how to determine when to act in the case that $t_{d}^{\operatornamewithlimits{e}}$ occurs before $t_{d}^{\operatornamewithlimits{o}}$.
@@ III-C Using Time as a Proxy for Exploration
We will assume accuracy increases monotonically with information gathered to rule out the impact of noise for our proof. For the purposes of discussing time as proxy for exploration, we will also assume information gain is passive.
@@
Assumption 5 is generally not the case, there is a large collection of research on interactive agents [sadigh2018planning, yang2018cm3, isele2019interactive, tian2021anytime, hu2023active]. But for the purposes of this section, we assume the ego vehicle’s behavior does not effect information gain to simplify the discussion on the importance of time.
Given the assumptions 3, 4 and 5, we will show that if $t_{d}^{\operatornamewithlimits{e}}$ occurs before $t_{d}^{\operatornamewithlimits{o}}$, then $t_{d}^{\operatornamewithlimits{e}}$ is the optimal time to make a decision under the catastrophic assertion. We will first analyze the case of two possible futures and then extend it to multiple possible futures.
@@
Suppose Assumptions 4 and 5 hold. A rational agent waiting until the last possible time to decide $t_{d}$ between two possible solution sets $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ can only reduce the chance of infinite possible loss, for an added cost that is strictly finite.
As a sketch of the proof we enumerate all possible cases of a choice made with less information and the corresponding rewards when the choice is either correct or incorrect. The full proof is in the Appendix [^2] Appendix at http://tiny.cc/ICRA2025 in Section -B. By Assumption 3, a best strategy delays a decision until $t_{d}$.
There can be multiple decisions and therefore multiple decision points over a given trajectory. Earlier decisions must be made first and often influence the ability to make later decisions. The following lemma concerns the uniqueness of state for a given decision point.
@@
Suppose Assumption 3 and 4 hold. A rational agent waiting until the last possible time to decide $t_{d}$ between multiple solution sets $\mathcal{T}_{i}\;\forall i\in\{1,...,m\}$ can only reduce the chance of infinite possible loss, for an added cost that is strictly finite.
@@
From Lemma 1 there is a unique decision point. At the time trajectory sets will split. These trajectory sets can be grouped, by common action, into meta-trajectory sets that split on the given decision. The proof of Theorem 1 concludes the proof. ∎
@@ III-D Maximum Entropy
To reach our final formulation for delayed decision making in the presence of multiple possible futures we return to Eqn. (3) and regularize it with the the maximum entropy formulation [ziebart2008maximum, eysenbach2019if]
@@ IV-A Quadratic Program Formulation
Given the piecewise linear bounds obtained from a cell planner [anon2024multi], we simultaneously solve for multiple trajectories corresponding to multiple futures where all trajectories are locked up to time $t_{d}$. The decision time $t_{d}$ could be found exactly by binary search, however, to reduce computation, an acceptable look ahead time is used as a tuneable parameter with a backup plan of the previous most probable solution in the event that the solver encounters an infeasibility.
@@
Figure 2: Validation in CARLA: (Left) Depiction of the scene in RVIZ. (Right) Diagram of the space-time graph showing the planned trajectory and corresponding velcoity. The ego vehicle approaches an intersection. The car currently crossing the intersection is depicted in gray on the ST graph. The car in the oncoming lane (shown in blue on the ST graph) is predicted as potentially turning, and potentially going straight. At 1s the trajectories are matched. After 1s, their are two trajectories corresponding to continuing straight and yielding for the turning car.
@@ IV-B Handling Complexity
The ST cell planner [anon2024multi] quickly finds all sets of constraints where each set of constraints corresponds to a different quadratic program. However, if there are $k$ traffic agents, in the worst case there may be $\bar{c}=2^{k}$ sets of quadratic constraints (quadratic programming problems), see Fig. 6 in Section -C of the Appendix for an example of this worst case complexity. This complexity is compounded when considering multiple futures.
There are $m$ joint predictions of the future, however many prediction algorithms only output single agent predictions. If each car has $r$ predictions, there will be $m=r^{k}$ possible futures resulting from the different future combinations for the each agent. Note, that if we were instead to consider the full action space of each agent, there would be an exponential number of possibilities being compounded at every time step. Additionally, when passing constraints for each future to the multi-future optimization, the selection of constraints is combinatoric with $\bar{c}^{m}$ possible problems as you try to optimize against the different sets of constraints for each future. This results in an $O(2^{k^{r^{k}}})$ number of optimization problems in the worst case, which is clearly computationally intractable. Here we discuss approximations that can be used to reduce the number of problems to a manageable amount. Section -F in the Appendix includes an example that helps illustrate the computational complexity of this section.

### v2

@@ I Introduction
Prediction technology continues to advance, and multiple prediction outputs are now a staple of state-of-the-art prediction methods [tang2019multiple, choi2019drogon, phan2020covernet, salzmann2020trajectron++, shi2022motion, isele2024gaussian]. This paper examines how an autonomous driving (AD) agent can utilize multiple predictions in the behavior planning process. In the context of this work, behavior planning corresponds to the combined task of decision-making and trajectory planning [slas].
@@
Consider the scenario depicted in Fig. 1. A pedestrian walks along a road and will likely continue straight (with 80% probability), but the pedestrian is positioned close to the street, indicating that they might turn to cross the street (with 20% probability). Selecting the most probable sequence of events results in an overly aggressive and risky behavior; we assume they will not cross and are wrong 20% of the time. However, a cautious policy unnecessarily brakes in the middle of an intersection for a pedestrian, creating confusing and potentially dangerous consequences. Intuitively, the solution in this instance is to try to balance the two behaviors: slow down so that it is possible (i) to brake if necessary, but (ii) not to make the overly conservative assumption that an unlikely event is true. This paper presents a formalism through which the problem of planning in the presence of multiple predictions can be expressed, where the intuitive solution of our example emerges as a result.
@@
Starting with probabilistic predictions, we can adopt a behavior that maximizes the expected reward [schmerling2018multimodal]. However, this is often insufficient for safety-critical applications because negative events, such as collisions, are discouraged but not forbidden. This shortcoming can be addressed by directly incorporating constraints into the formulation [ahn2021safe]. However, such an approach is still overly conservative; even though marginalizing over an agent’s distribution accurately assigns risk to both possible futures, it neglects that only one future can be true, generally resulting in a behavior that avoids both possibilities.
@@
Figure 1: Illustrative example: A pedestrian at the intersection has possible intentions of walking straight or turning right. The ego vehicle (at the center of the image) must comprehend that the pedestrian’s potential behaviors reflect that planning, e.g., by slowing down enough to stop safely if needed.
@@
One way around this difficulty is by delaying an agent’s decision while selecting an action that keeps a response to both futures possible [alsterda2019contingency, cui2021lookout]. In this work, we formalize the importance of delayed decision making and identify the optimal time to wait. While the generalized framework for behaving under the uncertainty associated with multiple predictions can be carried out by various planning methods, to ground the formalism, we specifically detail its application to an MPC framework. The special case of solving the problem with MPC has some similarity to existing work [alsterda2019contingency, chen2022interactive]. Our approach solves parallel optimizations which was also explored by Adajania et al. [adajania2022multi], though in a different problem setting. Similar to our approach, Contingency MPC [alsterda2019contingency] also solves for parallel plans with different costs and constraints, sharing a common initial state and action. However, they focus on steering under different friction assumptions, rather than handling predictions, and they issue a command for just a single time step. This avoids the decision process and the possibility of infeasibility. Branch MPC [chen2022interactive] takes into account another agent’s actions with corresponding probabilities, but unlike our work, Branch MPC reasons over the other agent’s entire action space which is substantially more computationally prohibitive, and therefore their investigation was limited to a single other agent.
@@
For an autonomous driving use case [tariq2022autonomous], there is a high non-convexity not addressed in many of the other works, which we show cannot be avoided in the worst case. Hence, we present reasonable approximations that allow us to handle the non-convexities efficiently. As part of the approximate solution, we develop a very fast piece-wise-linear trajectory planning approximation which can enable the quick evaluation of candidate solutions, and might be useful outside of our specific setting. Finally, to demonstrate the practicality of our algorithm, as well as the realtime computation speed, we showcase our results in simulation and on 1/10 scale cars.
@@ II-B Problem Formulation
The multi-future trajectory planning (MFTP) problem is posed as a stochastic game described by the tuple $\{\mathcal{S},\mathcal{A},P,\mathcal{R}\}$, where $\mathcal{S}$ is the world state which consists of a set of states $\{S^{1},\dots,S^{n}\}$ for $n$ agents including the ego agent, $\mathcal{A}=\{A^{1},\dots,A^{n}\}$ is the set of actions for agents, $P:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ is the set of transition probabilities, and =$\mathcal{R}=\{R^{1},\dots,R^{n}\}$ is the set of rewards for each agent. The sets of feasible states, actions, and reward function (1) for the ego agent are represented by $S^{e}$, $A^{e}$, and $R^{e}$, respectively. The ego agent generates a trajectory $\tau=\{s_{1},\dots,s_{T}\}$ consisting of a set of states where $s_{t}\in S^{e}$ and $T$ is the time horizon. A future $f=\{s_{1}^{j},\dots,s_{T}^{j}\},~{}\forall j\in[1,\dots,n]\backslash{e}$, is defined as the set of future states for all agents starting from the current time. This implies that the behavior of each non-ego agent is fixed over a future $f$.
@@
In practice, Assumption 1 allows us to make use of multi-intention prediction algorithms where the outputs can be taken directly from any off-the-shelf algorithm since there is no requirement that the predictions adhere to any particular constraints or have any particular distribution. Additionally, because we’re considering the trajectory as a whole, we remove the large computational burden of Branch MPC [chen2022interactive] since we no longer need to account for different actions from different agents at every time step.
@@
Note that it is sufficient to satisfy the constraints pertaining to the true future, however, that is unknown. Therefore, for $i\in\{1,\dots,m\}$, we define the corresponding belief of future $f_{i}$ being $f_{*}$:
@@ III Approach
We suppose that the probability of each future trajectory is invariant for each planning instance, which allows us to optimize the expected value:
$\max_{\tau}\mathbb{E}_{f\sim p(f)}[R(\tau|f)]=\max_{\tau}\sum_{i=1}^{m}p_{i}R(\tau|f_{i})\enspace.$ (3)
However, in the case of safety critical systems, some trajectories will be catastrophic. We acknowledge this by allowing for some trajectories to have a corresponding reward of $\infty$- ∞111in the case of autonomous driving, this will correspond to trajectories that result in collisions. This makes it difficult to work with expectations, since all expectations with non-zero probability of trajectories that result in $\infty$- ∞ reward will have $\infty$- ∞ expected value. We will therefore assert optimality can be sacrificed for the purpose of minimizing catastrophic events. We will refer to this assertion as the catastrophic assertion.
@@
Remark 2
Utilizing the worst-case formulation, our algorithm places utmost importance on maintaining safety and predictability. However, in the event that safety cannot be maintained, i.e., incurring a $\infty$- ∞ reward is inevitable, we can call upon our crash mitigation system to take evasive actions [tariq2023risk].
@@ III-A Maximum Probability Trajectory
To ensure we have a viable action in response to an uncertain future, we want to maximize the probability that our trajectory is in the feasible set $\mathcal{T}_{*}$ of the true future $f_{*}$ at the time $t_{R}$ that the true future is revealed. We are therefore interested in maximizing the joint probability that we satisfy constraint $c_{i}$ and $\mathcal{T}_{i}$ is in $\mathcal{T}_{*}$ for each future $i$ at time $t=t_{R}$. A convenience of working with trajectories is that we know that if the ego agent is on $\tau$ at time $t_{R}$ and $\tau\in\mathcal{T}_{*}$, then it can stay on a feasible trajectory for all time after $t_{R}$.
@@
In general, it is not possible to find one single trajectory that stays in all feasible spaces, as trajectories corresponding to different futures will inevitably diverge. However, we can maximize the number of constraints we are able to satisfy by keeping multiple constraints satisfied for as long as possible. Some readers might correctly identify the connection to maximum entropy here, which we will discuss in section III-D. However, we will first discuss how keeping many trajectories feasible is equivalent to delaying a decision.
@@ III-B Delayed Decisions
At some time $t=t_{R}$ we will know which future is the true future $f_{*}$. Trivially, this could be after time equal to the horizon time $T$ has passed, at which point the future will have already happened. In many cases, however, we can observe indications or definitive decisions in advance since certain actions make subsequent future actions impossible. Note that in order to make use of any observed indications or decisions, the ego agent needs sufficient time to react.
If we know $f_{i}=f_{*}$, we can determine which states keep us in $\mathcal{T}_{i}$. Without knowing $f_{*}$ in advance, the last moment $t_{d}^{\operatornamewithlimits{e}}$ the ego agent can make a decision is defined as the point in time where there is at least one feasible trajectory that belongs to all trajectory sets $\mathcal{T}_{1},\dots,\mathcal{T}_{m}$. Note that by symmetry, other agents have a similar $t_{d}^{\operatornamewithlimits{o}}$, after which point at least some futures will no longer be possible.
@@
Note that if, as our policy, we choose to make a conservative decision immediately before $t_{d}^{\operatornamewithlimits{e}}$, we get a risk-free reduction in uncertainty (and increase in safety) for all cases where $t_{d}^{\operatornamewithlimits{o}}$ happens first. Additionally, if $t_{d}^{\operatornamewithlimits{e}}$ happens earlier, we can still benefit from waiting as a result of additional acquired information. The next section describes how to determine when to act in the case that $t_{d}^{\operatornamewithlimits{e}}$ occurs before $t_{d}^{\operatornamewithlimits{o}}$.
@@ III-C Using Time as a Proxy for Exploration
We will assume accuracy increases monotonically with information gathered to rule out the impact of noise for our proof. For the purposes of discussing time as a proxy for exploration, we will also assume that information gain is passive.
@@
Assumption 4 is generally not the case; there is a large collection of research on interactive agents [sadigh2018planning, yang2018cm3, isele2019interactive, tian2021anytime, hu2023active]. But for the purposes of this section, we assume the ego vehicle’s behavior does not effect information gain to simplify the discussion on the importance of time.
Given the assumptions 2, 3 and 4, we will show that if $t_{d}^{\operatornamewithlimits{e}}$ occurs before $t_{d}^{\operatornamewithlimits{o}}$, then $t_{d}^{\operatornamewithlimits{e}}$ is the optimal time to make a decision under the catastrophic assertion. We will first analyze the case of two possible futures and then extend it to multiple possible futures.
@@
Suppose Assumptions 3 and 4 hold. A rational agent waiting until the last possible time to decide $t_{d}$ between two possible solution sets $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ can only reduce the chance of infinite possible loss, for an added cost that is strictly finite.
As a sketch of the proof we enumerate all possible cases of a choice made with less information and the corresponding rewards when the choice is either correct or incorrect. The full proof is in the Appendix [^2] Appendix at http://tiny.cc/ICRA2025 in Section -B. By Assumption 2, the best strategy delays a decision until $t_{d}$.
There can be multiple decisions and therefore multiple decision points over a given trajectory. Earlier decisions must be made first and often influence the ability to make later decisions. The following lemma concerns the uniqueness of the state for a given decision point.
@@
Suppose Assumption 2 and 3 hold. A rational agent waiting until the last possible time to decide $t_{d}$ between multiple solution sets $\mathcal{T}_{i}\;\forall i\in\{1,...,m\}$ can only reduce the chance of infinite possible loss, for an added cost that is strictly finite.
@@
From Lemma 1, there is a unique decision point. At the time, trajectory sets will split. These trajectory sets can be grouped, by common action, into meta-trajectory sets that split on the given decision. The proof of Theorem 1 concludes the proof. ∎
@@ III-D Maximum Entropy
To reach our final formulation for delayed decision making in the presence of multiple possible futures, we return to Eqn. (3) and regularize it with the the maximum entropy formulation [ziebart2008maximum, eysenbach2019if]
@@ IV-A Quadratic Program Formulation
Given the piecewise linear bounds obtained from a cell planner [anon2024multi], we simultaneously solve for multiple trajectories corresponding to multiple futures where all trajectories are locked up to time $t_{d}$. The decision time $t_{d}$ could be found exactly by binary search, however, to reduce computation, an acceptable look ahead time is used as a tunable parameter with a backup plan of the previous most probable solution in the event that the solver encounters an infeasibility.
@@
Figure 2: Validation in CARLA: (Left) Depiction of the scene in RVIZ. (Right) Diagram of the space-time graph showing the planned trajectory and corresponding velocity. The ego vehicle approaches an intersection. The car currently crossing the intersection is depicted in gray on the ST graph. The car in the oncoming lane (shown in blue on the ST graph) is predicted as potentially turning, and potentially going straight. At 1s, the trajectories are matched. After 1s, their are two trajectories corresponding to continuing straight and yielding for the turning car.
@@ IV-B Handling Complexity
The ST cell planner [anon2024multi] quickly finds all sets of constraints where each set of constraints corresponds to a different quadratic program. However, if there are $k$ traffic agents, in the worst case, there may be $\bar{c}=2^{k}$ sets of quadratic constraints (quadratic programming problems), see Fig. 6 in Section -C of the Appendix for an example of this worst case complexity. This complexity is compounded when considering multiple futures.
There are $m$ joint predictions of the future, however many prediction algorithms only output single agent predictions. If each car has $r$ predictions, there will be $m=r^{k}$ possible futures resulting from the different future combinations for each agent. Note, that if we were instead to consider the full action space of each agent, there would be an exponential number of possibilities being compounded at every time step. Additionally, when passing constraints for each future to the multi-future optimization, the selection of constraints is combinatoric with $\bar{c}^{m}$ possible problems as you try to optimize against the different sets of constraints for each future. This results in an $O(2^{k^{r^{k}}})$ number of optimization problems in the worst case, which is clearly computationally intractable. Here we discuss approximations that can be used to reduce the number of problems to a manageable amount. Section -F in the Appendix includes an example that helps illustrate the computational complexity of this section.